# Data-driven ODE modeling of the high-frequency complex dynamics of a fluid flow

### v1

@@ Abstract
Introduction. We are often eager to identify a dynamical system model that generates observable time series data. Many systems that describe physical phenomena have been found since the finding of the equation of motion. In decades, various methods have been proposed that estimate a dynamical system describing observed data combining physical knowledge and machine learning techniques. Chorin et al [1] proposed a method to describe unknown partial dynamics that are not described by equations of physical systems. Physics-informed machine learning methods [2, 3, 4] construct a model by supposing physical conditions that a system generating time series is expected to have. Sometimes, we do not have any physical knowledge of a system that generates time series. Several approaches have been proposed concerning modeling dynamics using machine learning from given time series data without physical knowledge. The dynamic model decomposition [5] is a method to derive macroscopic features and estimate a dynamical system of the variables. Reservoir computing [6, 7, 8, 9] is a recurrent neural network with low computational cost. These methods model a continuous dynamical system as discrete-time, not differential equations.
In some studies, a system of ordinary differential equations (ODEs) is constructed to describe the dynamics of deterministic time series. Neural ODE [10] recently attracted much attention for using a neural network to model ODE. The method based on the theory of the Koopman operator [11, 12, 13] projects the phase space to an infinite dimensional space to construct a model as a linear system. There are studies to make a simple and understandable model of ODEs in a finite dimensional space without a neural network. Several studies [14, 15, 16, 17] have estimated a system of ODEs from the time series of all variables of the background system. Other investigations [18, 19, 20] have derived a system of ODEs using an observable variable and its time derivatives as the model variables. In these studies, prior knowledge of the background dynamics was used to choose basis functions for the regression; however, this method is not appropriate for practical purposes.
@@
Table 1: Sets of parameters. These parameters are used for modeling the dynamics of the Rössler equation and fluid flow. 222 We set $(\delta_{grid},n,T,\Delta t,m,p)=(0.25,5\cdot 10^{4},10^{5},0.10,3,0.1)$ in the case of the Rössler equation, and $(\delta_{grid},n,T,\Delta t,m,p)=(0.40,5\cdot 10^{4},5\cdot 10^{4},0.05,3,0.1)$ in the case of the fluid flow, where the parameter $\delta_{grid}$ is the grid size used to determine the Gaussian radial basis function, $n$ is the number of regression data points, $T$ is the time length of the time series of the base and targeted variable, and $\Delta t$ is the time step of the time series data. The parameter $\sigma$ of $\phi_{j}$ is determined from $m$ and $p$ (see [22] for more details).
@@
The joint model reconstructs the connection between $E_{1}$ and $E_{10}$ (Fig. 6). This imitation of the connection between the two energy variables contributes to the success of describing the high-frequency variable, which is difficult to model. The positive Lyapunov exponents of the base model are $\lambda_{1}=0.2157,\lambda_{2}=0.0718,\lambda_{3}=0.0200$. We cannot obtain the actual Lyapunov exponents because the analytical derivation of a closed-form equation for the energy variable from the Navier–Stokes equations is an unsolved problem.

### v2

# Data-driven ODE modeling of the high-frequency complex dynamics via a low-frequency dynamics model

@@ Abstract
Introduction. We are often eager to identify a dynamical system model that generates observable time series data. Many systems that describe physical phenomena have been found since the finding of the equation of motion. In decades, various methods have been proposed that estimate a dynamical system describing observed data combining physical knowledge and machine learning techniques. Chorin et al [1] proposed a method to describe unknown partial dynamics that are not described by equations of physical systems. Physics-informed machine learning methods [2, 3, 4] construct a model by supposing physical conditions that a system generating time series is expected to have. Sometimes, we do not have any physical knowledge of a system that generates time series. Several approaches have been proposed concerning modeling dynamics using machine learning from given time series data without physical knowledge. The dynamic mode decomposition [5] is a method to derive macroscopic features and estimate a dynamical system of the variables. Reservoir computing [6, 7, 8, 9] is a recurrent neural network with low computational cost. These methods usually model a continuous dynamical system as discrete-time, not differential equations. Note that a model constructed using reservoir computing is commonly a discrete-time system, but there are some types of reservoir computing using differential equations to describe the dynamics of the reservoir states [12].
In some studies, a system of ordinary differential equations (ODEs) is constructed to describe the dynamics of deterministic time series. Neural ODE [13] recently attracted much attention for using a neural network to model ODE. The method based on the theory of the Koopman operator [14, 15, 16] projects the phase space to an infinite dimensional space to construct a model as a linear system. There are studies to make a simple and understandable model of ODEs in a finite dimensional space without a neural network. Several studies [17, 18, 19, 20] have estimated a system of ODEs from the time series of all variables of the background system. Other investigations [21, 22, 23] have derived a system of ODEs using an observable variable and its time derivatives as the model variables. In these studies, prior knowledge of the background dynamics was used to choose basis functions for the regression; however, this method is not appropriate for some practical purposes  [^1] In some situations, prior knowledge of the background dynamics can be obtained and the modeling with the knowledge is useful [10, 11]..
@@
Table 1: Sets of parameters. These parameters are used for modeling the dynamics of the Rössler equation and fluid flow. 333 We set $(\delta_{grid},n,T,\Delta t,m,p)=(0.25,5\cdot 10^{4},10^{5},0.10,3,0.1)$ in the case of the Rössler equation, and $(\delta_{grid},n,T,\Delta t,m,p)=(0.40,5\cdot 10^{4},5\cdot 10^{4},0.05,3,0.1)$ in the case of the fluid flow, where the parameter $\delta_{grid}$ is the grid size used to determine the Gaussian radial basis function, $n$ is the number of regression data points, $T$ is the time length of the time series of the base and targeted variable, and $\Delta t$ is the time step of the time series data. The parameter $\sigma$ of $\phi_{j}$ is determined from $m$ and $p$ (see [25] for more details).
@@
The joint model reconstructs the connection between $E_{1}$ and $E_{10}$ (Fig. 6). This imitation of the connection between the two energy variables contributes to the success of describing the high-frequency variable, which is difficult to model. The positive Lyapunov exponents of the base model are $\lambda_{1}=0.2157,\lambda_{2}=0.0718,\lambda_{3}=0.0200$, although we do not have the actual exponents for comparison. We cannot obtain the Lyapunov exponents using a governing equation, because a closed-form equation for the energy variable cannot be derived from the Navier–Stokes equations. Since the attractor dimension is high, it is hard to estimate Lyapunov exponents directly from time series [31].